Project

General

Profile

NDOS Commissioning Meeting 2011/01/06

Attendance: Jon, Denis, Mark Messier, Mat, Kurt, Dan, Athans, Alena, Mayly, Leon, Chad, Peter

Agenda:
  1. DCS Status & Plans
  2. DAQ/Electronics Update & Plans
  3. Shifters' report
  4. AOB

DCS:

  • Chad:
    • now writing env. data to the dB
    • will test more over the weekend, hope to have something to present at Monday's Comm. mtng.
  • Athans: made some progress recently, hope to have a demo of the sender <-> receiver apps sometime next week

DAQ:

  • Andrew:
    • we can read FEB status registers but:
    • have to switch to internal command mode to do that.
    • this switch disrupts data flow
    • requires a firmware fix, which is currently under testing.
  • Peter:
    • latest version of DCMProgUtils is now being tested; it appears just one file is out of synch. The development release will be updated tomorrow
    • MB triggers:
      • still no clear signals being found
      • Holger's code is doing the correct trigger logic
      • events should be ~304 us after our T0.
      • the search continues...
  • Mat:
    • new DSO data set collected yesterday, new hardware masks were generated; as a result, there was a net gain of ~500 channels.
    • some hardware work was done yesterday, where some APDs and optical connections were cleaned and I/V curves were taken. Details are in doc-dB 5492. Leon expressed concern about the procedure of the I/V curve data taking and warned of the potential of doing real damage to the APDs... the procedures will be worked out with Leon before any further data is collected.
    • Mat & Denis will remove ~10 more APDs from DCM09 (2-2).
  • The Plan:
    • work on DCM software and firmware improvements will continue
    • DSO data will be collected every Mon., Wed. and Fri. mornings, before work begins on the detector. Jon will communicate this to personnel.

Shifters' report: nothing specific, all goes reasonably well.